How would I find an equation of the line throught (2,2) that is parallel to the line passing through (4,5) and (10,-2)?
Answer by mananth(16946)

x1 y1 x2 y2
4 5 10 -2

slope m = (y2-y1)/(x2-x1)
( -2 - 5 )/( 10 - 4 )
( -7 / 6 )
m= -7/6
The equation of line parallel to the line passing through the above points will be the same = -7/6
m= -1 1/ 6

Plug value of the slope and point(2,2)
y=mx+b
2.00 = -7/6 * 2 +b
b= 2 + 7/3

b= 13/ 3
So the equation will be
Y = (-7/6) x + 13/3
m.ananth@hotmail.ca